The following table shows the number of cars sold last month by six dealers at Centreville Nissan dealership and their number of years of sales experience.

Years of Experience Sales
1 7
2 9
2 9
4 8
5 14
8 14

Management would like to use simple regression analysis to estimate monthly car sales using the number of years of sales experience. The standard error of the estimate is equal to ________.

Multiple Choice

  • 2.60

  • 5.00

  • 2.84

  • 1.84
Answer #1

Solution: The data given in the question statement is as below. Computing additional columns Years of Experience 1x 7.5 8.5 8SSE = Σ(y-97 = 13.50 SSE 13.50 n-2 6- 3.375 s= MSE =13.375 = 1.84 Hence, standard error of the estimate 1.84